When a number is subtracted from both the numerator and the denominator or 18/25 the result is 3/4.

Answer by oscargut(2103)

That number is -3 because
(18-(-3))/(25-(-3))= 21/28 = 3/4

You can ask me more at: mthman@gmail.com
Thanks